### v2

# EdiVal-Agent: An Object-Centric Framework for Automated, Fine-Grained Evaluation of Multi-Turn Editing

Abstract
Instruction-based image editing has advanced rapidly, yet reliable and interpretable evaluation remains a bottleneck. Current protocols either (i) depend on paired reference images—resulting in limited coverage and inheriting biases from prior generative models—or (ii) rely solely on zero-shot vision–language models (VLMs), whose prompt-based assessments of instruction following, content consistency, and visual quality are often imprecise.
To address this, we introduce EdiVal-Agent, an automated and fine-grained evaluation framework grounded in an object-centric perspective, designed to assess not only standard single-turn but also multi-turn instruction-based editing with precision. Given an input image, EdiVal-Agent first decomposes it into semantically meaningful objects, then synthesizes diverse, context-aware editing instructions while dynamically updating object pools across turns. These two stages enable two novel object-centric metrics tailored for multi-turn evaluation and one global metric of visual quality: 1) EdiVal-IF, which measures instruction following by combining open-vocabulary object detectors for symbolic checks with VLMs for semantic verification on detector-guided crops; 2) EdiVal-CC, which evaluates content consistency by calculating semantic similarity of unchanged objects and background using the evolving object pools; and 3) EdiVal-VQ, which quantifies changes in overall visual quality with human preference models.
Instantiating this pipeline, we build EdiVal-Bench, a multi-turn editing benchmark covering 9 instruction types and 13 state-of-the-art editing models spanning in-context  [^1] In this paper, we label certain closed-source models—GPT-Image-1, Nano Banana, and Gemini 2.0 Flash Image—as in-context, since they are integrated into autoregressive language models in the web UI and support in-context multi-turn editing., flow-matching, and diffusion paradigms. We further conduct experiments comparing multi-turn editing with single-shot complex editing, highlighting the distinctive characteristics of different model paradigms. We demonstrate that EdiVal-Agent can be used to identify existing failure modes, thereby informing the development of the next generation of editing models. Project page: https://tianyucodings.github.io/EdiVAL-page/
Figure 1: Overview of our workflow and representative model’s performance. For visualization, we adopt two thresholds: a consistency score of at least $90$ and a visual quality score of at least 66. Details of the automated evaluation pipeline are provided in Figure 2 and Section 2. In multi-turn editing, models exhibit distinct weaknesses: GPT-Image-1 struggles with content consistency, Qwen-Image-Edit underperforms in both visual quality and content consistency, and FLUX.1-Kontext-dev lags in instruction following, whereas Nano Banana shows no single dominant weakness.
1 Introduction
What truly defines the success of an image editor? At its core, editing requires making targeted, instruction-driven changes while preserving contextual consistency and perceptual realism—often across multiple refinement turns. Yet current evaluation practice struggles to capture this multifaceted objective.
When ground-truth edited images are available, a common strategy is to compare model outputs against these references (e.g., MagicBrush Zhang et al. (2023), UltraEdit Zhao et al. (2024), AnyEdit Yu et al. (2025), EmuEdit Sheynin et al. (2024)). Typical metrics include pixel-level distances (e.g., L1/L2) and semantic similarities (e.g., DINO Caron et al. (2021) and CLIP Radford et al. (2021)). While informative, such metrics suffer from two structural issues: (i) the space of acceptable edits is inherently large, whereas a single reference provides only one realization; and (ii) references are frequently synthesized by existing editing models (e.g., Prompt-to-Prompt Hertz et al. (2023), SDXL Podell et al. (2024), DALLE-2 Ramesh et al. (2022)), thereby importing their biases and limitations into the evaluation itself. Consequently, high reference similarity does not necessarily imply faithful instruction following, preservation of irrelevant content, or aesthetically plausible outcomes.
A complementary line of work employs zero-shot VLMs as interpretable evaluators (e.g., VIEScore Ku et al. (2023), GEdit-Bench Liu et al. (2025), I2EBench Ma et al. (2024), HQ-Edit Hui et al. (2024), Complex-Edit Yang et al. (2025), and ImgEdit Ye et al. (2025)) and queries VLMs about specific aspects of an edit. While VLMs offer holistic, language-mediated judgments, they remain insufficient for precise editing assessment for several reasons. First, for instruction-following evaluation, they are notoriously poor at spatial reasoning Zhang et al. (2025b); Chen et al. (2024); Chang et al. (2025) and are prone to object hallucinations in existence, category, attributes, and relations Bai et al. (2024). These issues together undermine their ability to assess common object-related edit instructions. Second, they have limited sensitivity to pixel-level changes and frequently miss subtle, localized modifications Vo et al. (2025) (e.g., fine structures, small attribute shifts, etc.), which are crucial for evaluating content consistency. Third, since they are predominantly pretrained on natural images rather than synthetic generations, their priors are miscalibrated for artifacts and aesthetics, leading to failures in detecting common generative defects (e.g., extra fingers) and in modeling perceptual “naturalness” Liang et al. (2024); Xu et al. (2023); Ma et al. (2025), which humans are sensitive to. Consequently, VLM-only scoring lacks the precision and reliability required for fine-grained evaluation across instruction following, content consistency, and visual quality. However, we find recently open-source state-of-the-art editing models (e.g., Qwen-Image-Edit Wu et al. (2025a), Step1X-Edit Liu et al. (2025), BagelDeng et al. (2025)) solely rely on VLMs for evaluation.
Figure 2: Framework of EdiVal-Agent. It first decomposes images into semantically meaningful objects, such as metal yellow sign and metal brown pole, and identifies their contextual relationships, e.g., they are both in foreground. It then generates diverse and proper editing scenarios at scale which are based on the initial analysis, e.g., Change the color of metal brown pole to gray. Finally, it systematically evaluates editing model outputs from multiple axes with our proposed metrics: EdiVal-IF, EdiVal-$C$, and EdiVal-VQ. Our agentic pipeline is agnostic to the expert tools used and can be readily enhanced with more advanced tools in the future.
To address these challenges, we introduce EdiVal-Agent: an automated and fine-grained evaluation agent for multi-turn instruction-based image editing from an object-centric perspective, designed to assess not only standard single-turn but also multi-turn instruction-based editing with precision. As shown in Fig. 2, EdiVal-Agent first decomposes it into semantically meaningful objects, then synthesizes diverse, context-aware editing instructions while dynamically updating object pools across turns. These two stages enable two novel object-centric metrics tailored for multi-turn evaluation and one global metric of visual quality: 1) EdiVal-IF, which measures instruction following by combining open-vocabulary object detectors for symbolic checks with VLMs for semantic verification on detector-guided crops; 2) EdiVal-CC, which evaluates content consistency by calculating semantic similarity of unchanged objects and background using the evolving object pools; and 3) EdiVal-VQ, which quantifies changes in overall visual quality with human preference models. We show that EdiVal-IF yields stronger agreement with human judgments in instruction-following evaluation compared to thresholded CLIP directional (CLIP_dir) scores Gal et al. (2022) and using VLMs alone, as evidenced in Sec. 2.5.
Instantiating the agentic pipeline, we curate a new multi-turn image editing benchmark, EdiVal-Bench, featuring 9 instruction types and 13 existing editing models—spanning in-context, flow-matching, and diffusion paradigms, across both closed- and open-source systems—conduct fine-grained analyses, and draw actionable insights. Empirically, as demonstrated in Fig. 1 and Tab. 3, GPT-Image-1 excels at instruction following yet ranks near the bottom in content consistency, whereas Seedream 4.0 and Nano Banana performs strongly on both axes. Besides, open-sourced models like Qwen-Image-Edit significantly degrade in instruction following and visual quality when editing turns increase, while FLUX.1-Kontext-max and FLUX.1-Kontext-dev lags in instruction following. We further contrasts multi-turn editing with single-shot complex prompts Yang et al. (2025), highlighting complementary strengths and failure modes. We hope that our agent pipeline, benchmark, and analyses accelerate the transition of multi-turn editing toward practical applications.
Table 1: Key attributes of open-source edit benchmarks. Note that ImgEdit Ye et al. (2025) does not include multi-turn editing experiments in the paper.
Benchmark # Size Object-centric Automated Multi-Turn Free from Ref. Images Tools used
EmuEditSheynin et al. (2024) 3,055 L1, CLIP, DINO
MagicBrushZhang et al. (2023) 1,053 L1, L2, DINO, CLIP
AnyEditYu et al. (2025) 1,250 L1, CLIP, DINO
I2EBenchMa et al. (2024) 2,240 VLM
GEdit-BenchLiu et al. (2025) 606 VLM
HQ-EditHui et al. (2024) 1,651 VLM
ImgEdit-BenchYe et al. (2025) 811 VLM
Kontext-BenchLabs et al. (2025) 1,026 Human Annotation
EdiVal-Bench (ours) 1,716 Detector, VLM, L1, DINO, HPS
Key contributions. 1) Agent: EdiVal-Agent is a fully automated evaluator that performs object-centric decomposition, generates diverse multi-turn editing instructions, and measures overall editing quality using two object-centric metrics (EdiVal-IF and EdiVal-CC) plus EdiVal-VQ for visual quality. 2) Benchmark: using EdiVal-Agent, we construct EdiVal-Bench with 1,716 instructions across nine types and three turns on 572 real-world images, with comparisons to prior benchmarks in Tab. 1. 3) Human agreement: EdiVal-IF attains 81.3% agreement with human ratings for instruction following, outperforming zero-shot VLMs and CLIP-based baselines. 4) Evaluation: we assess 13 editors (diffusion, flow-matching, and close source) along instruction following, content consistency, and visual quality. 5) Insights: overall ranking—Seedream 4.0 >> Nano Banana >> FLUX.1-Kontext-max >> GPT-Image-1; the strongest open-source editor, Qwen-Image-Edit, exhibits exposure bias under multi-turn editing. 6) Artifacts & settings: we reveal luminance drift across turns, and contrast multi-turn against complex single-shot editing to delineate strengths and weaknesses across model families.
2  EdiVal-Agent
2.1 Overview
As illustrated in Fig. 2. The pipeline comprises three stages: (1) Decomposition uses a VLM (e.g., GPT-4o; other VLMs are viable alternatives) to extract structured, object-level descriptions—objects, attributes, and relations—enabling symbolic reasoning; (2) Instruction Generation produces multi-turn, diverse, compositional prompts by maintaining an explicit object pool and sampling from nine instruction types spanning subject-, attribute-, relational-, text-, count-, and global-level edits; (3) Evaluation scores edited images with EdiVal-IF, Edival-CC, and EdiVal-VQ.
2.2 Step 1: Decomposition
Given an image, a VLM-based agent parses clearly visible foreground objects and returns per-object JSON with fields object, color, material, text, count, and a boolean foreground. Names follow "{material} {color} {object}"; unknown fields are omitted; person identity is never recorded (only wearables/accessories). Example: {"metal yellow sign": {"object":"sign","color":"yellow","material":"metal","text":"SCHOO
L","count":1,"foreground":true}}
. An aggregated all_objects string concisely lists objects (e.g., “metal yellow sign . metal brown pole”). We apply this stage to GEdit-Bench Liu et al. (2025) (606 images), exclude 34 images with sensitive personal content, and retain 572 images. After extraction, Grounding-DINO validates objects and detects bounding boxes; only reliable detections are kept to seed instruction generation and evaluation. The filtered objects are stored in the All_Objects_Pool and later used to initialize three distinct object pools that dynamically track the evolving state of instruction generation.
2.3 Step 2: Instruction Generation
From the decomposed scene, the agent generates multi-turn edits that are grounded in the current object state. The instruction taxonomy (nine types; six categories) appears in Table 2. We maintain three evolving pools at turn $t$: $\mathcal{P}^{\text{all}}_{t}$ (all objects ever present), $\mathcal{P}^{\text{unch}}_{t}$ (original objects not edited up to $t$), and $\mathcal{P}^{\text{avail}}_{t}$ (objects currently editable). With a turn budget MAX_TURNS, at each turn the agent (i) selects a type—defaulting to subject_add if $\mathcal{P}^{\text{avail}}_{t}=\emptyset$, otherwise sampling a type not yet used in the chain; (ii) selects object(s) from $\mathcal{P}^{\text{avail}}_{t}$; (iii) emits a natural-language instruction via GPT-4o referencing those objects and the scene state; and (iv) updates $\mathcal{P}^{\text{all}}_{t+1}$, $\mathcal{P}^{\text{avail}}_{t+1}$, and $\mathcal{P}^{\text{unch}}_{t+1}$ according to the intended edit. When a background_change edit applies at turn $t$, background-consistency scoring is disabled since this turn, and we append “make {objects_in_foreground} unchanged” to the instruction to preserve object-level comparability, where $\texttt{objects\_in\_foreground}=\{\,o\in P_{t}^{\mathrm{avail}}:\;o.\texttt{foreground}=\texttt{true}\,\}.$ The loop is adaptive by expanding/contracting $\mathcal{P}^{\text{avail}}_{t}$ and naturally compositional. Our default sets MAX_TURNS=3=3 (In our implementation, each turn is assigned a distinct instruction type.), though longer chains are easily obtained by allowing repetition or adding types.
Table 2: Instruction types in EdiVal-Bench created by EdiVal-Agent, grouped by semantic category. Counts are shown per turn (T1–T3).
Category Instruction Type Example Instruction T1 T2 T3 Total
Subject-centric subject_add Add bench on the left of metal red fire hydrant. 67 77 93 237
subject_remove Remove wooden brown door. 75 69 61 205
subject_replace Replace stone gray railing with wooden fence. 54 57 55 166
Attribute-centric color_alter Change the color of metal white airplane to blue. 56 73 57 186
material_alter Change the material of plastic black pen to metal. 66 50 72 188
Text-related text_change Replace the text ’BEARS CONTROL’ on cotton black cap with ’WILD PATH’. 64 70 54 188
Relational position_change Change the position of ceramic white cup to right of plastic white laptop. 52 63 48 163
Counting count_change Change the count of fur brown bear to 3. 73 58 60 191
Global background_change Change the background to forest, remain the brown fur bear unchanged. 65 55 72 192
2.4 Step 3: Evaluation
The first two stages enable two novel object-centric metrics for multi-turn editing evaluation for instruction following and content consistency, respectively, and one global metric for visual quality:
EdiVal-IF
To evaluate instruction following, we introduce EdiVal-IF, which assesses multi-turn edits by comparing the image from the previous turn, $I^{t-1}$, to the current image, $I^{t}$. For a given instruction $P^{t}$ at turn $t$, the score is determined differently for symbolically and semantically verifiable tasks. Symbolically verifiable types ($T_{\text{sym}}$)—such as subject_add, subject_remove, subject_replace, position_change, and count_change—are evaluated using an open-vocabulary object detector $\mathcal{M}_{\text{detect}}$ Liu et al. (2024b). The detector’s outputs, including bounding boxes and confidence, are assessed against geometric and logical criteria $\mathcal{F}_{\text{sym}}$ derived from the instruction. For example, for a position_change instruction “Move [A] to the left of [B]”, $\mathcal{F}_{\text{sym}}$ verifies that the x-coordinate of A’s bounding box $\mathcal{B}$ center is less than that of B in $I^{t}$, i.e., $\text{center}_{x}(\mathcal{B}_{A}^{t})<\text{center}_{x}(\mathcal{B}_{B}^{t})$. In this case,
$\text{EdiVal-}{\text{IF}}(I^{t},I^{t-1},P^{t}\in T_{\text{sym}})=\mathcal{F}_{\text{sym}}(\mathcal{M}_{\text{detect}}(I^{t-1},I^{t}|P^{t})).$ (1)
Semantically verifiable types ($T_{\text{sem}}$)—color_alter, material_alter, text_change, and background_change—are evaluated with a VLM $\mathcal{M}_{\text{VLM}}$ Yang et al. (2024). To focus the evaluation, the VLM is applied to detector-guided object crops ($I_{\text{o}}$) using instruction-specific templates.
$\text{EdiVal-}{\text{IF}}(I^{t},I^{t-1},P^{t}\in T_{\text{sem}})=\mathcal{M}_{\text{VLM}}(I^{t-1}_{\text{o}},I^{t}_{\text{o}}|P^{t})=\mathcal{M}_{\text{VLM}}(\mathcal{M}_{\text{detect}}(I^{t-1},I^{t}|P^{t})).$ (2)
We show that EdiVal-IF achieves superior human agreement (Sec. 2.5). The multi-turn editing success rate is defined as the logical AND of the EdiVal-IF scores across all edits along the chain, whereas the marginal task rate at turn $t$ is defined according to the formulas 1 and 2 provided above.
(a) Base image
(b) GPT-Image-1
(c) FLUX.1-max
Figure 3: Beautification vs. preservation under the prompt: “Change the background to a library.” GPT-Image-1 tends to increase HPSv3 via beautification, while FLUX.1-Kontext-max emphasizes fidelity to the input.
EdiVal-$C$
To assess content consistency, EdiVal-CC measures the preservation of non-target content between the base image $I^{0}$ and the current image $I^{t}$. Given editing instructions $P^{1:t}$ from turn 1 to turn $t$, the object pools $\mathcal{P}^{\text{unch}}_{t}$ and $\mathcal{P}^{\text{all}}_{t}$ are dynamically updated. Let $\Omega$ denote the entire image area. Using object bounding boxes from the base image ($\mathcal{B}^{0}_{o}$) and the current image ($\mathcal{B}^{t}_{o}$), extracted by the detector $\mathcal{M}_{\text{detect}}$, the background region is defined as $\Omega_{\text{bg}}^{t}=\Omega-\bigcup_{o\in\mathcal{P}^{\text{all}}_{t}}(\mathcal{B}^{0}_{o}\cup\mathcal{B}^{t}_{o})$, i.e., the region obtained by excluding all objects that have appeared. Background consistency is then computed as $s^{t}_{\text{bg}}=\phi(I^{0}_{\text{bg}},I^{t}_{\text{bg}})$, where $I^{t}_{\text{bg}}=\Omega_{\text{bg}}^{t}\circ I^{t}$ denotes the background of the image, and $\phi$ is a similarity function such as $L_{1}$ distance or DINO-based similarity. For unchanged objects, we compute the per-object consistency $s^{t}_{o}=\phi(I^{0}_{o},I^{t}_{o})$ for each $o\in\mathcal{P}^{\text{unch}}_{t}$, and then average them. Formally, the final EdiVal-CC score emphasizes semantic preservation by averaging the feature-level similarities of the background and unchanged objects (see Appendix. I.3 for details):
$\text{EdiVal-CC}(I^{t},I^{0},P^{1:t})=\tfrac{1}{2}\left(s_{\text{bg}}^{t}+\tfrac{1}{|\mathcal{P}^{\text{unch}}_{t}|}\sum_{o\in\mathcal{P}^{\text{unch}}_{t}}s^{t}_{o}\right).$ (3)
EdiVal-CC aligns with the intuitive notion of consistency, providing a precise measurement.
EdiVal-VQ.
Zero-shot VLMs are not trained for reliable assessment of image quality—particularly artifacts and aesthetics—and we find they are imprecise as scoring functions (see Appx. H). Consequently, we adopt Human Preference Score v3 (HPSv3) Ma et al. (2025) as our visual-quality (VQ) evaluator. In practice, applying preference models to unedited, real photographs often yields relatively low aesthetic scores. We also observe divergent behaviors across editors (See Fig. 3): some (e.g., GPT-Image-1) tend to beautify images and increase HPSv3, whereas others (e.g., FLUX.1-Kontext-max) preserve the original appearance with minimal aesthetic change. Because aesthetic preference is inherently user- and task-dependent, and beautification can trade off with content consistency (already incorporated into our overall score), we report EdiVal-VQ separately and do not fold it into the aggregate metric.
EdiVal-O.
We aggregate Instruction Following (EdiVal-IF) and Content Consistency (EdiVal-CC) into a single overall score. Since both metrics are unit-free and normalized to $[0,1]$ but capture complementary aspects, we follow prior work and use the geometric mean to balance them and penalize imbalance (Liu et al., 2025; Ku et al., 2023). Formally, $\text{EdiVal-O}\;=\;\sqrt{\text{EdiVal-IF}\;\cdot\;\text{EdiVal-CC}}\,.$
Design Scope and Limitations.
We omit style_change from our taxonomy because style categories are inherently ill-defined, which makes instruction-following (EdiVal-IF) difficult to evaluate reliably. Extending EdiVal-Agent with style-aware recognition is promising future work. After language-based extraction, we validate objects using Grounding-DINO Liu et al. (2024a) and prune low-confidence or ambiguous detections. This stabilizes the object pool and reduces error propagation during instruction generation and IF evaluation. By default, we employ Grounding-DINO as the open-vocabulary detector, Qwen2.5-VL as the VLM, and DINOv3 Siméoni et al. (2025) as the image feature extractor due to their state-of-the-art performance and open-source availability, which facilitates community use. The agentic pipeline is tool-agnostic and can be readily strengthened by substituting more advanced modules in the future.
2.5 Measuring Human Agreement
Figure 4: Results of human agreement. Dashed lines represent the average accuracy of each method. EdiVal-IF achieves 81.3% human agreement accuracy, significantly outperforming the VLM (Qwen2.5-VL) at 75.2% and thresholded CLIP_dir at 65.4%. Note that the CLIP_dir threshold is tuned separately for each task.
Setup.
We conduct human study on edits made by four exemplary models, Step1X-Edit, AnyEdit, Gemini 2.0 Flash and Flux.1-Kontext-dev, on EdiVal-Bench, generated by EdiVal-Agent as described in Sec. 2.3. In total, we collect 4,5764{,}576 human annotations of edits. During evaluation, raters were shown the original image, the edited image, and the corresponding instruction, and asked a binary question: “Evaluate whether the edited image successfully follows the given instruction.”
Results.
Figure 4 summarizes the findings. EdiVal-IF achieves a human agreement accuracy of 81.3%, significantly higher than VLM-only (Qwen-2.5-VL, 75.2%), CLIP_dir (65.4%), and other zero-shot VLMs. These results verify that integrating VLMs reasoning with object detection leads to better alignment with human judgment compared to existing methods. The inter-annotator’s agreement rate (85.5%) indicates the best performance any evaluation tool can reach.
We attribute the improvement in instruction-following evaluation to two factors. First, for symbolically verifiable instruction types—subject_add, subject_remove, subject_replace, position_change, and count_change—EdiVal-IF relies solely on Grounding-DINO. It determines the success of an edit by checking object presence/absence, the positions of object centers, and the number of bounding boxes. Results for position_change and subject_remove show that these fixed rules, combined with Grounding-DINO, can significantly outperform Qwen2.5-VL in edit evaluation. We hypothesize that errors in position_change stem from poor spatial reasoning, while failures in subject_remove are due to hallucinations regarding object existence. Second, semantically verifiable types—color_alter, material_alter, text_change, and background_change—are evaluated using Qwen2.5-VL combined with Grounding-DINO. The decomposition stage in EdiVal-Agent can supports evaluation by localizing text regions, enabling the LLM to reason more precisely about text edits. These findings indicate that EdiVal-IF not only enhances interpretability but also improves the practical applicability of evaluation pipelines in real-world settings that demand human-like understanding. Nonetheless, EdiVal-IF has failure modes, which we document and analyze in Appendix. E.
3 Benchmarking Multi-turn Editing
Table 3: Results of multi-turn editing. EdiVal-IF, EdiVal-$C$ , and EdiVal-O across three sequential editing turns. Best per column in dark red; second-best in lighter red.
Latency EdiVal-IF EdiVal-$C$ EdiVal-O Rank
Technique Model In-Context Date (s/img) T1 T2 T3 T1 T2 T3 T1 T2 T3
Unknown Seedream 4.0 25.09.10 15.78 75.93 55.58 41.59 92.51 88.03 85.86 83.81 69.95 59.76 1
Nano Banana 25.08.26 9.70 70.70 50.66 35.35 93.91 90.48 89.48 81.48 67.70 56.24 2
GPT-Image-1 25.07.16 71.30 73.12 54.89 38.35 81.00 77.78 75.50 76.96 65.34 53.81 3
Gemini 2.0 Flash 25.02.05 8.20 68.07 45.96 28.42 90.58 85.10 80.88 78.52 62.54 47.94 5
Flow Matching FLUX.1-Kontext-max 25.06.03 10.34 69.49 46.89 31.83 93.93 90.90 88.40 80.79 65.29 53.04 4
Qwen-Image-Edit 25.08.04 115.08 72.90 44.06 22.55 84.22 80.52 77.98 78.36 59.56 41.93 6
Step1X-Edit 25.04.25 20.42 61.89 34.97 17.83 92.76 88.52 85.21 75.77 55.64 38.98 7
FLUX.1-Kontext-dev 25.06.25 29.21 59.97 32.69 16.61 95.32 92.24 90.22 75.61 54.91 38.71 8
OmniGen 24.09.11 19.70 54.72 24.48 10.66 93.00 88.42 83.92 71.34 46.52 29.91 9
Diffusion AnyEdit 24.11.24 3.93 41.07 16.32 7.22 86.42 78.91 70.10 59.58 35.89 22.50 10
UltraEdit 24.07.07 3.15 51.37 17.70 6.36 86.80 84.50 82.40 66.78 38.67 22.89 11
MagicBrush 23.06.16 4.08 42.31 15.73 4.90 86.96 81.26 76.86 60.66 35.75 19.41 12
IP2P 23.12.15 4.09 37.41 10.66 2.80 76.85 68.36 60.30 53.62 26.99 12.99 13
Summary of Results.
Figure 5: Marginal Task Success rate across turns.
Table 3 shows that Seedream 4.0 achieves the strongest overall performance, leading in EdiVal-O across all three turns (83.81/69.95/59.76) with competitive latency (15.8 s/img) [^2] Closed-source latencies are measured in the provider’s hosted web UI; open-source latencies on a single NVIDIA A100 GPU with default settings.. Nano Banana offers the best speed–quality trade-off at 9.7 s/img, ranking second in EdiVal-O and staying close to Seedream 4.0 in both instruction following and consistency. GPT-Image-1 excels in instruction following, but its very high latency (71.3 s/img) and weaker consistency lower its overall score, consistent with more regenerative behavior that prioritizes aesthetics over stability. Among open-source systems, Qwen-Image-Edit performs well initially (EdiVal-O 78.36 at T1) but degrades rapidly with additional turns, likely due to exposure bias as discussed below. We can see that there is a clear gap between the performance of closed-source and open-source systems. With the exception of Qwen-Image-Edit, our model rankings exactly match those reported on the Artificial Analysis leaderboard (rank by human vote) as of September 12, 2025; see Appendix. D.
3.1 Instruction Following
Marginal Task Success Rate.
For a given turn, the marginal task success rate (Eqns. 1 and 2) is the proportion of prompts for which the edit requested at that turn is successfully executed. By contrast, the instruction-following score in Table 3 reports the multi-turn task success rate at turn $i$: the logical AND of the EdiVal-IF scores across all edits along the chain. Figure 5 summarizes per-turn performance. High-ranking models—such as Seedream 4.0, Nano Banana, and FLUX.1-Kontext-max—maintain relatively stable EdiVal-IF across turns, even though Seedream 4.0 and FLUX.1-Kontext-max are not in-context editors (they do not condition on prior prompts or intermediate images). In contrast, several other models exhibit a clear decline in marginal success as the number of turns increases.
A particularly salient case is Qwen-Image-Edit. Although it is the strongest open-source system at turn 1 (EdiVal-O $78.36$ vs. $81.48$ for Nano Banana), its performance degrades more rapidly over subsequent turns. We hypothesize that this stems from exposure bias (Ning et al., 2023; Schmidt, 2019): many single-turn editors are trained to operate on real images and ground-truth inputs rather than on their own previous outputs. When asked to edit their own generations, small distributional mismatches compound across turns, reducing stability; this effect is further aggravated when the model can only attend to a limited history.
Marginal Task Success Rate Across Instruction Types.
We analyze marginal subtask success rates across turns for different instruction types. The results for Nano Banana are shown in Fig. 7. Other editing models exhibit similar behavior. Nano Banana performs relatively well on attribute-centric tasks such as color_alter and material_alter, but poorly on position_change and count_change, indicating weaknesses in spatial and numerical reasoning, respectively.
Figure 6: Marginal task success rate grouped by task types for Nano Banana.
Figure 7: Illustration of background consistency. Instruction: “Remove beige brick house.” The grounding box is the union of all object regions from the raw and edited images.
3.2 Content Consistency
We evaluate two aspects: (i) unchanged-object consistency (Fig. 8), which measures whether objects that are not edited up to turn $i$ remain unchanged, and (ii) background consistency (Fig. 7), which assesses whether the background remains stable when it is not explicitly modified. When calculating consistency, the grounding box is extracted from the raw input image and applied to all edited images. We therefore choose to report DINOv3 over $L_{1}$ distance for consistency computation because even small shifts in object location can lead to large variations in pixel-wise $L_{1}$ loss, even if unchanged objects are well preserved. By relying on DINO features, we ensure that consistency is measured semantically, capturing attributes such as object identity, attributes, and texture, etc. Nevertheless, the consistency scores from DINOv3 remain highly correlated with those computed using pixel-wise $L_{1}$ loss (See results in the Appendix. J). Based on the results, the most consistent editing model is FLUX.1-Kontext-dev, followed by Nano Banana and FLUX.1-Kontext-max. In contrast, GPT-Image-1 ranks near the bottom, showing notably poor consistency across turns. Representative qualitative examples are shown in Figure 8 and Figure 7.
(a) Base Image
(b) GPT-Image-1 (95.19)
(c) Nano Banana (98.05)
(d) Qwen-Image-Edit (94.96)
Figure 8: Illustration of object consistency. Instruction: “Remove brick beige house.” The grounding box, extracted from the raw input image, highlights the localized region used to compute unchanged-object consistency. The corresponding consistency score is shown in brackets.
3.3 Visual Quality
Besdies EdiVal-VQ, we report the absolute change in VQ score relative to the base image: $\text{EdiVal-VQ}\Delta_{i}\;=\;\bigl|\text{EdiVal-VQ}_{\text{turn }i}-\text{EdiVal-VQ}_{\text{base}}\bigr|.$ Smaller $\Delta$ indicates stronger style fidelity to the base image; larger $\Delta$ reflects greater beautification or stylistic drift. As summarized in Table 4, GPT-Image-1 achieves the highest aesthetic scores across turns and the biggest $\Delta$, indicating a significant stylistic shift (Fig. 3). For preserving the base image’s look (small $\Delta$), Gemini 2.0 Flash shows the least drift, with Nano Banana also performing well. We provide low-level exposure statistics analysis in Appendix. G.
[…]
Technique Model HPS $\Delta$
T1 T2 T3 T1 T2 T3
Unknown Seedream 4.0 5.14 5.15 5.15 0.76 0.77 0.77
Nano Banana 4.94 5.12 5.26 0.56 0.73 0.88
GPT-Image-1 6.65 6.59 6.56 2.27 2.21 2.18
Gemini 2.0 Flash 4.44 4.23 4.07 0.05 0.15 0.32
Flow Matching FLUX.1-Kontext-max 5.12 5.07 5.04 0.41 0.49 0.47
Qwen-Image-Edit 5.86 5.72 5.15 1.47 1.34 0.77
Step1X-Edit 4.06 3.34 2.76 0.33 1.04 1.63
FLUX.1-Kontext-dev 5.12 5.07 5.04 0.73 0.69 0.65
OmniGen 4.61 4.07 3.50 0.23 0.31 0.89
Diffusion AnyEdit 3.66 2.80 1.95 0.72 1.58 2.44
UltraEdit 4.79 4.68 4.36 0.41 0.30 0.02
MagicBrush 3.85 3.08 2.36 0.54 1.30 2.02
IP2P 3.20 2.38 1.44 1.18 2.01 2.94
Figure 9: Turn-3 instruction following: Multi-turn vs. single-shot complex prompts.
3.4 Multi-turn Editing vs. Complex Editing
We compare two strategies for composing multiple edits. In multi-turn editing, instructions are executed sequentially—apply instruction 1, then apply instruction 2 to the result, and so on. In complex editing, we concatenate $C$ instructions into a single prompt and perform one edit (“complex level” $C$, with $C\in\{1,2,3\}$). Empirically (Fig. 9), when a model does not suffer from exposure bias, multi-turn editing tends to yield higher success rates, consistent with a step-by-step “chain of edits” (analogous to chain-of-thought in reasoning). For instance, Nano Banana benefits from the multi-turn formulation. Conversely, when exposure bias is pronounced, compressing instructions into a single, complex prompt can perform better; see Qwen-Image-Edit in Fig. 9.
4 Conclusion
We introduced EdiVal-Agent, an automated, and interpretable framework for evaluating instruction-based image editing. By leveraging symbolic object decomposition, structured instruction generation, and a hybrid evaluation pipeline integrating both specialist tools and vision-language reasoning models, EdiVal-Agent enables fine-grained, object-centric assessment of modern editing systems. Our design emphasizes transparency, extensibility, and real-world applicability—providing a new standard for evaluating multi-turn, compositional visual editing. We hope EdiVal-Agent will serve as a valuable resource for benchmarking, diagnosing, and advancing the next generation of instruction-based editing models. Limitations and discussions are deferred to Appendix B.
Ethics Statement
Our work focuses on developing reliable and interpretable evaluation methods for instruction-based image editing. While such technology holds promise for creative design, accessibility, and efficient content creation, it may also be misused for harmful purposes such as generating misleading, deceptive, or inappropriate content. We emphasize that our benchmark and evaluation framework are intended solely for advancing research in safe and trustworthy generative AI. To mitigate risks, we build on publicly available datasets, apply safety filters to generated images, and encourage responsible use aligned with ethical standards and community guidelines.
Reproducibility statement
We provide complete prompting templates and pseudo-code in the Appendix, along with implementation details and API links. Comprehensive results, datasets, and evaluation metrics are also documented. To ensure full reproducibility, we will release all code, data, and model checkpoints upon acceptance of this manuscript.
Appendix A Related Work
Instruction-based editing models.
InstructPix2Pix (IP2P) Brooks et al. (2023) introduced a two-stage recipe that converts a text-to-image diffusion model Rombach et al. (2022); Zhang et al. (2025a) into an editor: (i) synthesize paired editing data using Stable Diffusion Rombach et al. (2022) and training-free techniques such as Prompt-to-Prompt Hertz et al. (2023); (ii) fine-tune the diffusion model on these pairs. Subsequent systems—MagicBrush Zhang et al. (2023), UltraEdit Zhao et al. (2024), and AnyEdit Yu et al. (2025)—scale this paradigm to large, fine-grained real-image editing. More recent work (e.g., OmniGen Xiao et al. (2025); Wu et al. (2025b), Step1X-Edit Liu et al. (2025), FLUX.1 Kontext Labs et al. (2025), and Qwen-Image-Edit Wu et al. (2025a), Seedream Gao et al. (2025)) adopts task-aware architectures and increasingly leverages flow matching Liu et al. (2022); Lipman et al. (2022); Zhang et al. (2024).
A complementary line explores autoregressive (AR) editors such as Gemini 2.0 Flash Image Gemini2 (2025), Gemini 2.5 Flash Image (“Nano Banana”) Deepmind (2025), and GPT-Image-1 OpenAI (2025). These models enable in-context multi-turn editing: users iteratively refine an image within a conversational interface, with the model maintaining a coherent editing history. To our knowledge, we provide the first systematic comparison of in-context multi-turn AR editing versus context-free multi-turn editing with non-AR models across instruction following, content consistency, and visual quality.
Editing evaluation.
Early evaluations (e.g., Brooks et al. (2023)) rely on CLIP-based similarity Radford et al. (2021), including directional variants Gal et al. (2022), to approximate editing quality. However, CLIP emphasizes semantic alignment and is less sensitive to fine, pixel-level changes. When ground-truth edited images exist, it is natural to compare model outputs against references using pixel distances ($L_{1}$) and semantic similarities (DINO Caron et al. (2021), CLIP Radford et al. (2021)) Zhang et al. (2023); Zhao et al. (2024); Yu et al. (2025); Sheynin et al. (2024). Yet references are imperfect: the space of valid edits is inherently multimodal, while a single reference captures only one realization; moreover, many references are themselves synthesized by prior editors (e.g., Prompt-to-Prompt Hertz et al. (2023), SDXL Podell et al. (2024), DALLE-2 Ramesh et al. (2022)), importing their biases into evaluation.
Recent work relies exclusively on VLMs as interpretable judges—e.g., VIEScore Ku et al. (2023), HQ-Edit Hui et al. (2024), and Complex-Edit Yang et al. (2025)—by querying models such as GPT-4o OpenAI (2025) about specific aspects of an edit. While VLMs offer holistic, language-mediated assessments, they are insufficient on their own: they are notoriously poor at spatial reasoning Zhang et al. (2025b); Cheng et al. (2024); Chen et al. (2024); Qharabagh et al. (2024); Chang et al. (2025) and are prone to object hallucinations in existence, category, attributes, and relations Bai et al. (2024); they have limited sensitivity to pixel-level changes and frequently miss subtle, localized modifications Vo et al. (2025) (e.g., fine structures, small attribute shifts, etc.), which are crucial for evaluating content consistency; they are miscalibrated for artifacts and aesthetics Liang et al. (2024); Xu et al. (2023); Ma et al. (2025), which humans are sensitive to. Our approach, EdiVal-Agent, addresses these gaps by integrating VLM-based reasoning with grounding tools, symbolic, object-centric pixel- and semantic-level tools, and human preference models, yielding a precise and interpretable evaluation of instruction-based editing.
Editing tasks.
We consider three settings: (i) Single-turn vs. multi-turn. Multi-turn editing Zhang et al. (2023); Zhao et al. (2024) is more demanding than single-turn, as the model must maintain consistency across sequential instructions. In contrast to context-free multi-turn pipelines (each turn consumes the previous image and the next instruction), AR models Gemini2 (2025); Deepmind (2025); OpenAI (2025) support in-context multi-turn editing by conditioning on the full conversational history. (ii) Complex single-shot vs. multi-turn. Following Yang et al. (2025), a sequence of edits can be concatenated into a single complex prompt and executed in one pass; we compare this setting to genuine multi-turn editing. (iii) Other tasks. We focus on instruction-based editing, the most common regime; other scenarios (e.g., prompt-to-prompt/caption-to-caption Hertz et al. (2023)) are outside our scope. To the best of our knowledge, this paper offers the first comprehensive comparison covering single-turn, multi-turn, and complex single-shot editing within a unified framework.
Appendix B Limitation and Discussion
Given the object-centric evaluations conducted in this study, several limitations warrant consideration. First, our instruction types are limited to object-centric prompts, which may not capture the full range of creative editing requests typical in real-world scenarios. Future research should explore a broader spectrum of instructions, including those involving stylistic changes or complex narrative elements. Additionally, while our work provides a reliable and comprehensive evaluation framework for multi-turn editing, it does not apply the evaluation results to improve the editing models themselves. A straightforward extension would be to use evaluation scores for Best-of-N selection to improve inference-time performance. Future work could also explore post-training methods such as reinforcement learning, treating the evaluation scores as reward signals.
Appendix C More metrics for human agreement
The most straightforward metric is accuracy. Here, we provide more metrics measuring human agreement: Pearson Linear Correlation Coefficient (PLCC), Cohen’s Kappa Coefficient (Kappa) and F1 scores as shown in Tab. 5. However, we note that for 0/1 predictions, correlation metrics like PLCC and Kappa may be considered not suitable for measuring the agreement with human annotators.
Table 5: Task-specific and overall performance comparison across models. Metrics reported: Pearson Linear Correlation Coefficient (PLCC), Cohen’s Kappa, and F1. Best per column highlighted in bold.
Task Type Model PLCC Kappa F1
subject_add CLIP_dir 0.2110 0.1764 0.6914
Qwen2.5-VL 0.5331 0.5264 0.7893
EdiVal-IF 0.5365 0.5364 0.7786
background_change CLIP_dir -0.0329 -0.0076 0.8745
Qwen2.5-VL 0.5792 0.5686 0.8889
EdiVal-IF 0.5244 0.5157 0.8763
subject_remove CLIP_dir -0.0011 -0.0002 0.6592
Qwen2.5-VL 0.1891 0.1758 0.4896
EdiVal-IF 0.5473 0.5409 0.7837
count_change CLIP_dir 0.0456 0.0142 0.0782
Qwen2.5-VL 0.1998 0.1748 0.2162
EdiVal-IF 0.3431 0.3274 0.3571
material_alter CLIP_dir 0.2086 0.2038 0.4561
Qwen2.5-VL 0.8658 0.8616 0.8971
EdiVal-IF 0.4778 0.4624 0.6364
color_alter CLIP_dir 0.1841 0.0874 0.8542
Qwen2.5-VL 0.8409 0.8407 0.9573
EdiVal-IF 0.7820 0.7744 0.9338
position_change CLIP_dir 0.0996 0.0430 0.3285
Qwen2.5-VL -0.0381 -0.0374 0.1798
EdiVal-IF 0.3907 0.3271 0.5000
text_change CLIP_dir 0.6178 0.6173 0.8063
Qwen2.5-VL 0.7161 0.6947 0.8651
EdiVal-IF 0.7438 0.7347 0.8571
subject_replace CLIP_dir 0.0420 0.0121 0.8219
Qwen2.5-VL 0.6028 0.5994 0.8699
EdiVal-IF 0.5533 0.5429 0.8410
Overall CLIP_dir 0.3186 0.2568 0.6858
Qwen2.5-VL 0.6162 0.6161 0.7922
EdiVal-IF 0.6278 0.6273 0.8030
Appendix D Artificial Analysis Leaderboard
We report the leaderboard from the Artificial Analysis website as of September 12, 2025 (Fig. 10). To ensure a fair comparison, we align on the intersection of models evaluated by both platforms and exclude Qwen-Image-Edit. Among the overlapping systems—Seedream 4.0, Nano Banana (Gemini 2.5 Flash), GPT-Image-1 (GPT-4o), FLUX.1-Kontext-max, and Gemini 2.0 Flash—the relative ordering of human votes on Artificial Analysis matches our EdiVal rankings exactly (Table 6), supporting the accuracy of our methodology.
Table 6: Model rankings on the overlapping set. Relative ranks from Artificial Analysis (human votes) vs. EdiVal (ours) as of Sep 12, 2025.
Model Artificial Analysis (Rank) EdiVal (Rank)
Seedream 4.0 1 1
Nano Banana (Gemini 2.5 Flash) 2 2
GPT-Image-1 (GPT-4o) 3 3
FLUX.1-Kontext-max 4 4
Gemini 2.0 Flash 5 5
Figure 10: Artificial Analysis leaderboard (Sep 12, 2025). Screenshot of the public leaderboard used for comparison in Table 6.
Appendix E Failure Case
We discuss a representative failure mode of our evaluation. The most severe errors arise from false positives in Grounding-DINO, despite its strong open-vocabulary performance. Consider the prompt: “Replace [stone gray railing] with [wooden fence].” As shown in Fig. 11, Seedream 4.0 produces an edit that is visually correct. Our rule for subject_replace declares success if, on the edited image, the source object (stone gray railing) is no longer detected while the target object (wooden fence) is detected. However, Grounding-DINO occasionally reports both the source and target objects in the same region with high confidence, incorrectly suggesting that the source object remains and thereby degrading the measured instruction-following accuracy. Improving the reliability of open-vocabulary detection—particularly reducing false positives—would directly improve the fidelity of our evaluation.
(a) Base image
(b) Seedream 4.0 (edited)
(c) Detected: stone gray railing
(d) Detected: wooden fence
Figure 11: Failure due to detector false positives. Although the edit visually replaces the railing with a wooden fence, Grounding-DINO fires on both “stone gray railing” and “wooden fence” in overlapping regions, causing an incorrect failure in our instruction-following metric.
Appendix F Discussion on single-shot complex editing
Figure 12 shows that marginal success for the final instruction remains largely stable as complex prompt length increases. Together with the multi-turn drops seen in Figure 5, this pattern supports an exposure-bias explanation: performance degradation primarily stems from error accumulation across sequential edits rather than an intrinsic inability to handle multiple instructions in a single prompt.
Appendix G Discussion on Visual Quality
Beyond instruction following and content consistency, the perceptual quality of the edited image is a key dimension. We therefore report (i) a learned aesthetic score and (ii) several low-level image statistics that can surface systematic artifacts and drift in multi-turn editing pipelines.
Table 7: Turn-3 instruction following: Multi-turn vs. single-shot complex prompts, grouped by technique. Bold indicates which setting is higher for each model.
Technique Model Multi-turn (T3) Complex (C3)
In-Context Nano Banana 35.35 28.14
GPT-Image-1 38.35 28.78
Gemini 2.0 Flash 28.42 21.89
Flow Matching Qwen-Image-Edit 22.55 27.62
Step1X-Edit 17.83 15.73
FLUX.1-Kontext-dev 16.61 19.58
OmniGen 10.66 11.01
Diffusion AnyEdit 7.22 2.80
UltraEdit 6.36 8.22
MagicBrush 4.90 4.55
IP2P 2.80 2.80
Figure 12: Marginal task-success rate of the last instruction as a function of complex prompt length (levels $C=1,2,3$).
Figure 13: Per-image 99.9% luminance quantile across turns. Higher values indicate more extreme bright pixels and greater risk of over-exposure.
Low-level image statistics
In addition to learned aesthetic scores, we compute several low-level image statistics that help reveal systematic, multi-turn editing artifacts. Concretely, we convert RGB pixels to luminance using the Rec. 709 luma coefficients: Y= 0.2126R+ 0.7152G+ 0.0722B,Y\;=\;0.2126\,R\;+\;0.7152\,G\;+\;0.0722\,B, and for each edited image we extract the 99.9% luminance quantile (the per-image pixel value below which 99.9% of pixels fall). The 99.9% quantile is sensitive to high-exposure pixels and therefore highlights over-exposure and bright streaks while being robust to single-pixel outliers. In Figure 13 we plot the trend of this statistic across turns.
The measured trend shows a clear pattern: Qwen-Image-Edit and several other flow-matching models (with the notable exception of FLUX.1-Kontext-dev) exhibit a pronounced increase in the 99.9% luminance quantile over turns, indicating progressive brightening and increased risk of over-exposure. By contrast, regeneration-style editors such as GPT-Image-1 tend to produce lower luminance values than the input (reflecting darker, more conservative reconstructions), and several models remain stable across turns.
Figure 14 provides qualitative examples from Qwen-Image-Edit. The edited images exhibit elevated luminance and noticeable high-frequency bright artifacts (e.g., white streaks or “line” textures) that degrade perceptual quality, with luminance quintiles increasing substantially. Correspondingly, HPS drops from 6.19 to 4.19 and 3.34, suggesting that HPS is sensitive to over-exposure to some extent. In contrast, when querying VLMs about the visual quality of these images, the returned scores do not change in the first two turns and remain consistently above 50, reflecting a positive evaluation under the [0, 100] scale, while the T2/T3 edited images show significant artifacts.
(a) HPS: 4.25
VLM: 85
Luminance:0.7
(b) HPS: 6.19
VLM: 85
Luminance: 0.60
(c) HPS: 4.19
VLM: 85
Luminance: 0.97
(d) HPS: 3.34
VLM: 60
Luminance: 1.00
Figure 14: Representative Qwen-Image-Edit examples illustrating over-exposure and bright artifact formation across turns. Although editing instructions are often satisfied, the images show elevated luminance and high-frequency bright streaks that accompany the edits (visible especially in T2/T3). Editing instructions: [Remove polyester white skirt, Change the count of tennis ball to 4, Change the color of tank top to blue]. Note that VLM gives a positive score to all the images.
Appendix H VLMs failing to judge visual quality
The following is the zero-shot prompt for visual quality with VLMs. The example results are shown in Fig. 14.
You are an expert at evaluating image visual quality and naturalness.
I will show you an image.
Please analyze whether the image is visually pleasing and natural. Consider:
1. Is the image visually pleasing?
2. Is the image natural?
3. Does the image look natural and coherent?
Respond only with a score between 0 and 100, where 100 is the highest score.
100 means the image is visually pleasing and natural.
0 means the image is not visually pleasing and natural.
50 means the image is neutral.
Appendix I Algorithmic Details
This appendix provides the algorithmic details of our pipeline: object discovery and grounding-based filtering (decomposition), instruction generation for multi-turn editing, and evaluation (instruction following, consistency, and perceptual quality). We also list the exact prompts and implementation specifics needed for reproducibility, and summarize the model-generation configurations.
I.1 Decomposition
We first enumerate visible objects in an input image using a vision-language model (VLM) prompt, then filter these objects via visual grounding.
  • Object listing: We use GPT-4o with the prompt in Section I.2.1. The model returns a JSON with one entry per object and a terminal aggregated string key ‘‘All Objects’’.
  • Grounding filter: We use GroundingDINO SwinT-OGC Liu et al. (2024a) to retain only objects that can be visually grounded. We resize images to $512\times 512$. We keep detections meeting text/box thresholds (0.35) and reject oversized boxes by checking width/height in normalized coordinates; we use max_box_size=0.9 and filter large regions if area >0.4>0.4. The output augments each kept object with grounding counts, phrases, boxes, and scores, and creates a ‘‘Filtered All Objects’’ string listing retained objects.
Algorithm 1 Object Listing and Grounding Filter
1:image II
2:object pool $\mathcal{O}$ with grounding metadata
[…]
5:for all $(\text{name},\text{attrs})\in J$ excluding key All Objects do
6:  $(\text{boxes},\text{phrases},\text{scores})\leftarrow\textsc{Ground}(I,\text{name})$ $\triangleright$ thresholds $0.3\text{--}0.4$
7:  if $\text{boxes}\neq\emptyset$ and each box has $w,h<0.9$ and $\mathrm{area}\leq 0.4$ then
8:   $\mathcal{O}[\text{name}]\leftarrow\text{attrs}$; attach grounding metadata (count, boxes, phrases, scores)
9:  end if
10:end for
[…]
I.2 Instruction Generation
We generate multi-turn editing instructions from the grounded object pool. We support nine task types: local edits {subject_replace, subject_remove, material_alter, color_alter, subject_add, text_change, position_change, count_change} and the global edit {background_change}. We set MAX_TURNS=3. At each turn, we sample a new task type without repetition where feasible. Feasibility is checked against the current object pool (e.g., position_change requires at least two objects). If a sampled task is infeasible, we fall back to subject_add. We maintain an available-objects pool that is updated after each instruction according to its semantics (adds, removes, or modifies attributes). If a background change occurs, we mark bg_consistency=false for subsequent turns and restrict the pool to foreground objects for the remainder of the episode.
Algorithm 2 Multi-Turn Instruction Generation
1:grounded pool $\mathcal{O}_{0}$, turns T=3T{=}3
2:tasks $\{\tau_{t}\}$, instructions $\{I_{t}\}$, formats $\{F_{t}\}$, flag has_bg, set all_objects_ever
[…]
5:for $t=1$ to $T$ do
6:  $\text{cand}\leftarrow\{\text{all task types}\}\setminus\text{used}$
7:  $\tau_{t}\leftarrow\text{sample}(\text{cand})$;
8:  if not feasible($\tau_{t},\mathcal{O}$) then $\tau_{t}\leftarrow\texttt{subject\_add}$
9:  end if
10:  $F_{t}\leftarrow\text{format\_instruction}(\tau_{t},\mathcal{O})$ $\triangleright$ Query VLM by prompts in Section I.2.2
11:  $I_{t}\leftarrow\text{render\_instruction}(F_{t},\tau_{t},\mathcal{O})$ $\triangleright$ strip brackets; add unchanged list for background
12:  $\text{used}\leftarrow\text{used}\cup\{\tau_{t}\}$; append $F_{t},I_{t}$
13:  Update all_object_pool by adding any objects introduced in instruction $I_{t}$.
14:  Update available_object_pool by adding or removing objects as specified in $I_{t}$.
15:  Update unchanged_objects_pool by removing any objects affected by $I_{t}$.
16:  if $\tau_{t}=\texttt{background\_change}$ then
17:   $\text{has\_bg}\leftarrow\text{true}$; $\mathcal{O}\leftarrow\text{filter\_foreground}(\mathcal{O})$
18:  end if
19:end for
20:return $\{\tau_{t}\},\{I_{t}\},\{F_{t}\},\text{has\_bg},\text{all\_objects\_ever}$
Prompts (Full Text)
Below we reproduce the prompts used by our generators, reformatted for readability in print (content preserved).
I.2.1 Object Listing Prompt
You will be given an image. Your task is to identify and describe all clearly visible objects in the image in a structured JSON format.
Output rules:
  1. 1.
    Each object must be listed as a key in the JSON, using the format: “{material} {color} {object name}”. If the material or color is unknown, omit that part. Do not include any visible text in the key. Do not use “person” as an object name; instead, describe wearable items (e.g., “blue cotton shirt”).
  2. 2.
    For each object, the value is a dictionary with fields: “object” (type, e.g., shirt, cup), “color” (dominant color, use null if unknown), “material” (likely material, use null if unknown), “text” (visible text, null if none), “count” (number of instances), and “foreground” (boolean).
  3. 3.
    Do not include objects that are too small to describe, mostly occluded/incomplete, or only background scenery (e.g., distant sky, wall, floor).
  4. 4.
    Add a final key “All Objects” whose value is a single string listing all object names, formatted as: “{material} {color} {object name}. {color} {object name}. {material} {object name}. {object name}.” Exclude “null”/“None” and separate entries by “. ” (period + space). Do not include any text content in this list.
Example output (abridged JSON):
  • “cotton blue shirt”: {object: “shirt”, color: “blue”, material: “cotton”, text: null, count: 1, foreground: true}
  • “ceramic white cup”: {object: “cup”, color: “white”, material: “ceramic”, text: “GOOD DAY”, count: 1, foreground: false}
  • “leather bag”: {object: “bag”, color: null, material: “leather”, text: null, count: 2, foreground: true}
  • “red scarf”: {object: “scarf”, color: “red”, material: null, text: null, count: 1, foreground: true}
  • “All Objects”: “cotton blue shirt. ceramic white cup. leather bag. red scarf.”
I.2.2 Task Prompts
Subject Replace
You are given an image and asked to suggest a replacement object for a specific object in the scene.
Given object to replace: object_name
Your task:
  1. 1.
    Understand the scene context.
  2. 2.
    Suggest a new object that naturally replaces “object_name”.
  3. 3.
    Ensure the suggestion is realistic for the scene.
  4. 4.
    Respond with only the object name (e.g., “chair”, “lamp”, “book”).
Examples: In a kitchen: “bowl”, “mug”; on a street: “bus”, “truck”; in an office: “stool”, “bench”.
Answer format: New object name:
Material Alter
You are given an image and asked to suggest a new material for a specific object.
Object: object_nameCurrent material: current_material
Your task:
  1. 1.
    Identify the object.
  2. 2.
    Suggest a realistic alternative material that is easy to distinguish from the current one.
  3. 3.
    Respond with only the material name (e.g., “wood”, “metal”, “plastic”, “leather”).
Examples: cup: ceramic, glass, metal, plastic; chair: wood, metal, plastic, fabric; bag: leather, canvas, nylon, fabric.
Answer format: New material:
Position Change
You are given an image and asked to create a position change instruction.
Available objects: available objectsPositions: left, right, above, below
Your task:
  1. 1.
    Select a target object to move and a reference object.
  2. 2.
    Choose a relative position (left, right, above, below).
  3. 3.
    Ensure the instruction is physically reasonable.
  4. 4.
    Format: “Change the position of [target object] to [position] of [reference object]”.
Examples: “Change the position of [cup] to [right] of [book]”; “Change the position of [lamp] to [above] of [table]”.
Answer format: Position change instruction:
Count Change
You are given an image and asked to create a count change instruction.
Available objects: available objectsTarget count: target count
Your task:
  1. 1.
    Identify a suitable object for the requested count.
  2. 2.
    Ensure the target count is realistic for the scene.
  3. 3.
    Format: “Change the count of [object name] to [target count]”.
Examples: “Change the count of [cup] to [3]”; “Change the count of [book] to [2]”.
Answer format: Count change instruction:
Text Change
You are given an image and asked to generate new text content.
Context: text situation
Your task:
  1. 1.
    Generate text that fits the scene.
  2. 2.
    Keep text short: max 2 words in English or 4 Chinese characters.
  3. 3.
    Respond with only the text content (no quotes or extra words).
Examples: coffee shop: “COFFEE”, “OPEN”; book: “NOVEL”, “GUIDE”; sign: “EXIT”, “STOP”; Chinese: “ 咖啡 ”, “ 出口 ”.
Answer format: New text:
Color Alter
You are given an image and asked to suggest a new color for a specific object.
Object: object nameCurrent color: current color
Your task:
  1. 1.
    Suggest a simple, common color that fits the object.
  2. 2.
    Use only basic colors: red, blue, green, yellow, black, white, brown, gray, orange, purple, pink.
  3. 3.
    Choose a color different from the current color and answer with the color name only.
Answer format: New color:
Subject Add
You are given an image and asked to suggest a new object to add to the scene.
Reference object: reference objectPosition: position
Your task:
  1. 1.
    Propose an object that would naturally fit at the specified position relative to the reference object.
  2. 2.
    Ensure the suggestion is realistic and contextually appropriate.
  3. 3.
    Respond with only the object name (e.g., “lamp”, “book”, “cup”).
Examples: next to a desk: “chair”, “lamp”, “computer”; near a kitchen counter: “bowl”, “plate”, “mug”; by a window: “plant”, “curtain”, “book”.
Answer format: New object:
Background Change
You are given an image and asked to suggest a new background for the scene. The existing objects should remain unchanged.
Your task:
  1. 1.
    Propose a new background that works with the current setting.
  2. 2.
    Keep it simple and realistic; use 1–2 words (e.g., “kitchen”, “office”, “garden”, “beach”, “forest”).
  3. 3.
    Respond with only the background name.
Answer format: New background:
I.3 Evaluation
We evaluate in two modes: (i) Multi-turn (each turn edits the output of the previous turn), and (ii) Complex Editing (compress all instructions to a single prompt).
Instruction Following.
We compute a binary success per instruction with a detector combining GroundingDINO Liu et al. (2024a) and a VLM (Qwen2-VL-7B) Bai et al. (2025). Representative details:
  • Detector thresholds. Unless noted per task, GroundingDINO thresholds are 0.3–0.4; detections return normalized boxes $x_{1},y_{1},x_{2},y_{2}]$.
  • Cropping and small objects. For object-level checks we crop by detected boxes; very small boxes (<0.05<\!0.05 in width/height) can be enlarged before VLM queries.
  • Replace. Detect old and new objects in source/target; success if both are detected and any IoU between a source box (old) and a target box (new) is >0>0. A VLM pre-check rejects obvious non-replacements. See details in Alg 3.
  • Remove. Detect the object in the source; success if the object is absent in the target. See details in Alg 4.
  • Position change. Detect target and reference objects and verify the requested spatial relation using object centers; also ensure the object count did not increase spuriously. See details in Alg 6.
  • Count change. Use the detector to locate instances of the target object and take the number of validated detections as the count. See details in Alg 7.
  • Color/material. Crop the object in the target and ask the VLM a yes/no question about the new color/material. See details in Alg 8 and Alg 9.
  • Text change. If the instruction adds text anywhere, run the VLM on the whole image; if it replaces text on a specific object, first crop that object’s box, ask the VLM to extract the text, and compare it to the requested text. See details in Alg 10.
  • Background change. Ask the VLM yes/no whether the requested background category is present. See details in Alg 11.
Consistency.
We measure object and background stability as follows:
  • Object consistency (unchanged objects): DINOv3 ViT-B/16 Siméoni et al. (2025) feature similarity between crops of unchanged objects in base vs. target; we also report pixel L1 consistency and average across objects per image.
  • Background consistency: detect objects in all_objects_pool in base/target (GroundingDINO), mask them to isolate background, then compute masked L1 between backgrounds (optionally DINOv3 masked similarity). Background consistency is evaluated only when no background change occurred earlier (bg_consistency=true).
Perceptual Quality.
We report HPSv3 Ma et al. (2025) plausibility and aesthetics, plus luminance metrics. Quality is not folded into the Overall score.
Algorithm 3 Evaluate Subject Replace
1:base $B$, target $T$, old object name $o$, new object name $n$
2:success flag succ
[…]
4:if $S\neq\emptyset\ \wedge\ T_{n}\neq\emptyset$ then
5:  $\text{succ}\leftarrow\max_{b\in S,\,t\in T_{n}}\textsc{IoU}(b,t)>0$
6:else
7:  $\text{succ}\leftarrow\text{false}$
8:end if
9:return succ
Algorithm 4 Evaluate Subject Remove
1:base $B$, target $T$, object name $o$
2:success flag succ
[…]
5:return succ
Algorithm 5 Evaluate Subject Add
1:base $B$, target $T$, new object name $n$, optional reference object name $r$, optional position $p\in\{\text{left},\text{right},\text{above},\text{below}\}$
2:success flag
[…]
4:if $T_{n}=\emptyset\ \vee\ B_{n}\neq\emptyset$ then
5:  return false
6:end if
7:if $r$ and $p$ are provided then
8:  $B_{r}\leftarrow\textsc{Detect}(B,r,\tau)$;  $T_{r}\leftarrow\textsc{Detect}(T,r,\tau)$
9:  if $T_{r}=\emptyset$ then
10:   return false
11:  end if
12:  Choose max logits boxes $t\in T_{n}$, $u\in T_{r}$
13:  $(x_{t},y_{t})\leftarrow\textsc{Center}(t)$;  $(x_{u},y_{u})\leftarrow\textsc{Center}(u)$
14:  if $p=\text{left}\wedge x_{t}<x_{u}-\varepsilon_{x}$ then
15:   return true
16:  end if
17:  if $p=\text{right}\wedge x_{t}>x_{u}+\varepsilon_{x}$ then
18:   return true
19:  end if
20:  if $p=\text{above}\wedge y_{t}<y_{u}-\varepsilon_{y}$ then
21:   return true
22:  end if
23:  if $p=\text{below}\wedge y_{t}>y_{u}+\varepsilon_{y}$ then
24:   return true
25:  end if
26:  return false
27:else
28:  return true
29:end if
Algorithm 6 Evaluate Position Change
1:base $B$, target $T$, target object name $a$, reference object $r$, position $p$
2:success flag
[…]
5:if $T_{a}=\emptyset\ \vee\ T_{r}=\emptyset$ then
6:  return false
7:end if
8:if $|T_{a}|>|B_{a}|$ then
9:  return false $\triangleright$ No count inflation
10:end if
11:Select max logits boxes $t\in T_{a}$, $u\in T_{r}$
[…]
13:if $p=\text{left}$ then
14:  return $x_{t}<x_{u}-\varepsilon_{x}$
15:end if
16:if $p=\text{right}$ then
17:  return $x_{t}>x_{u}+\varepsilon_{x}$
18:end if
19:if $p=\text{above}$ then
20:  return $y_{t}<y_{u}-\varepsilon_{y}$
21:end if
22:if $p=\text{below}$ then
23:  return $y_{t}>y_{u}+\varepsilon_{y}$
24:end if
25:return false
Algorithm 7 Evaluate Count Change
1:target $T$, name $o$, requested count $c^{*}$
2:success flag
[…]
4:return $(\hat{c}=c^{*})$
Algorithm 8 Evaluate Color Alter
1:target image $T$, object name $o$, color $k$
2:return $\textsc{VLMYesNo}(T,\text{``Is the }o\text{ }k\text{?''})$
Algorithm 9 Evaluate Material Alter
1:target image $T$, object name $o$, material $m$
2:success flag
3:return $\textsc{VLMYesNo}(T,\text{``Is the }o\text{ made of }m\text{?''})$
Algorithm 10 Evaluate Text Change
1:target $T$, desired text $t^{*}$ (optionally object name)
2:success flag
[…]
4:Normalize $t$ and $t^{*}$ (case, punctuation, whitespace)
5:return $\textsc{text-match}(t,t^{*})$
Algorithm 11 Evaluate Background Change
1:target $T$, category $g$
2:success flag
3:return $\textsc{VLMYesNo}(T,\text{``Does the background show }g\text{?''})$
I.4 Overall Score and Aggregation Details
Let $\alpha_{t}$ be the image success rate at turn $t$: the fraction of images for which all edits up to and including turn $t$ are successful (aggregated per task type, then averaged). Let $\kappa$ denote the average content-consistency score combining object and background DINOv3 similarities when applicable.
  • Overall score. We report
    $\text{Overall}=\bigl[\operatorname{mean}_{t}(\alpha_{t})\;\times\;\operatorname{mean}(\kappa)\bigr]^{1/2}.$
  • Missing outputs across turns. For summary tables, we include only images that produce all required outputs for the evaluated mode. If a model fails to generate a later turn, that image is omitted from later-turn aggregates for that mode. Some edits will be rejected by some models since the sensitive content flag.
  • No unchanged objects. If the unchanged-object list is empty, object consistency is recorded as None and excluded from averages; background consistency is still computed when bg_consistency=true.
  • Turn-level reporting. We also report per-turn (T1, T2, T3) instruction-following and consistency, and per-task-type success rates $\alpha_{t,\text{type}}$. Quality metrics are reported separately and are not folded into Overall.
I.5 Model generations
We evaluate a mix of closed- and open-source editors using each model’s default settings (no hyperparameter tuning):
Modes.
For clarity in the paper: we report both Multipass and Complex Editing (renamed from singlepass for consistency with the rest of the paper).
Reproducibility Notes.
Prompts are provided in full (Section I.2); thresholds are specified above. Grounding uses SwinT-OGC weights; consistency uses DINOv3 ViT-B/16; the quality head follows our RAHF implementation, and HPSv3 is included when available. All other parameters are left at defaults.
Appendix J Additional Evaluation Results
In this section, we provide extended evaluation results. We separate the analysis into two modes: multi-turn editing and complex editing. Each mode is evaluated across three aspects: instruction following, consistency, and quality.
For the multi-turn editing mode, the overall instruction-following success rate is reported in Table 8, while success rates for individual instruction types appear in Tables 9 and 10. Consistency results are summarized in Table 14. We also observed that some input images are non-square after resizing, which can leave black padding on the top/bottom or left/right edges. Certain editing models, such as GPT-Image-1 and Qwen-Image-Edit, attempt to fill these areas, whereas others preserve them. To account for this, we separately report consistency for square (Table 15) and non-square inputs (Table 16). The conclusions remain consistent with the overall evaluation. Quality results for multi-turn editing are presented in Table 18.
For the complex editing mode, the overall instruction-following success rate is shown in Table 11, and per-instruction-type results are in Tables 12 and 13. Consistency and quality results are reported in Tables 17 and 19, respectively.
In consistency table, p99 means $99\%$ quantile of luminance value, and p999 means $99.9\%$ quantile of luminance value.
Table 8: Image success rates and overall task success rates across turns. (Multi-turn model)
Model Image Success Rate Overall Task Rate
T1 T2 T3 T1 T2 T3
Seedream 4.0 75.93 55.58 41.59 75.93 75.58 76.11
Nano Banana 70.70 50.66 35.35 70.70 72.59 68.24
GPT-Image-1 73.12 54.89 38.35 73.12 74.44 73.12
FLUX.1-Kontext-max 69.49 46.89 31.83 69.49 69.11 70.43
Gemini 2.0 Flash 68.07 45.96 28.42 68.07 67.72 68.42
Qwen-Image-Edit 72.90 44.06 22.55 72.90 62.94 56.12
Step1X-Edit 61.89 34.97 17.83 61.89 59.09 53.32
FLUX.1-Kontext-dev 59.97 32.69 16.61 59.97 56.29 51.40
OmniGen 54.72 24.48 10.66 54.72 48.60 42.48
UltraEdit 51.37 17.70 6.36 50.52 36.54 31.47
AnyEdit 41.07 16.32 7.22 40.03 39.34 40.56
MagicBrush 42.31 15.73 4.90 42.31 40.73 41.26
IP2P 37.41 10.66 2.80 37.41 32.87 34.27
Table 9: Task success rates (%) across five instruction types and three turns (multi-turn mode).
Model Subject Replace Subject Remove Material Alter Color Alter Subject Add
T1 T2 T3 T1 T2 T3 T1 T2 T3 T1 T2 T3 T1 T2 T3
Seedream 4.0 90.74 91.23 88.89 68.92 47.69 50.00 95.31 96.00 95.77 100.00 98.59 100.00 83.08 89.61 81.52
Nano Banana 91.84 92.31 75.47 64.18 51.61 40.35 91.94 89.36 87.50 100.00 97.18 98.11 73.02 77.94 72.41
GPT-Image-1 84.31 94.64 85.71 70.77 55.93 47.37 96.83 95.65 87.88 100.00 97.06 100.00 80.95 72.46 72.41
FLUX.1-Kontext-max 92.31 88.89 88.00 67.16 55.56 56.36 87.30 79.07 80.30 100.00 95.65 98.04 77.05 71.23 72.73
Gemini 2.0 Flash 83.33 92.98 78.18 58.67 53.62 50.82 90.91 82.00 83.33 100.00 89.04 98.21 77.61 72.73 75.27
Qwen-Image-Edit 87.04 82.46 70.91 70.67 31.88 37.70 93.94 90.00 79.17 100.00 97.26 94.74 77.61 55.84 39.78
Step1X-Edit 90.74 96.49 67.27 53.33 30.43 21.31 95.45 80.00 87.50 100.00 100.00 91.23 64.18 57.14 45.16
FLUX.1-Kontext-dev 85.19 80.70 72.73 54.67 42.03 32.79 84.85 74.00 73.61 100.00 98.63 94.74 67.16 61.04 39.78
OmniGen 88.89 84.21 58.18 46.67 21.74 19.67 84.85 72.00 70.83 100.00 90.41 91.23 53.73 51.95 37.63
UltraEdit 88.89 63.16 38.18 26.67 5.80 6.56 87.88 66.00 63.89 98.21 80.82 78.95 38.81 23.38 9.68
AnyEdit 74.07 66.67 61.82 37.33 39.13 36.07 78.79 68.00 68.06 78.57 68.49 78.95 22.39 38.96 25.81
MagicBrush 83.33 75.44 63.64 28.00 18.84 18.03 83.33 86.00 80.56 94.64 87.67 91.23 37.31 41.56 37.63
IP2P 75.93 66.67 65.45 25.33 8.70 18.03 74.24 70.00 65.28 87.50 82.19 75.44 23.88 28.57 19.35
Table 10: Task success rates (%) across four instrcution types and three turns (multi-turn mode).
Model Text Change Position Change Count Change Background Change
T1 T2 T3 T1 T2 T3 T1 T2 T3 T1 T2 T3
Seedream 4.0 95.31 97.14 96.23 39.22 39.68 48.94 18.06 19.30 20.00 98.46 96.36 93.06
Nano Banana 83.33 86.36 81.25 20.00 47.37 45.24 23.19 9.62 10.91 96.67 94.44 90.00
GPT-Image-1 88.33 97.01 97.96 31.11 40.68 50.00 11.27 18.52 18.18 98.39 94.44 91.30
FLUX.1-Kontext-max 80.36 86.15 82.69 18.75 38.18 44.44 8.82 9.09 12.28 96.88 92.59 92.54
Gemini 2.0 Flash 90.32 94.29 96.30 21.15 28.57 31.91 10.96 7.14 10.00 86.15 83.64 80.56
Qwen-Image-Edit 98.44 92.86 72.22 21.15 34.92 33.33 12.33 1.72 0.00 98.46 80.00 77.78
Step1X-Edit 60.94 51.43 44.44 13.46 31.75 27.08 0.00 1.72 1.67 87.69 87.27 83.33
FLUX.1-Kontext-dev 50.00 41.43 27.78 15.38 26.98 33.33 0.00 1.72 5.00 90.77 80.00 77.78
OmniGen 29.69 35.71 18.52 17.31 22.22 20.83 5.48 5.17 0.00 76.92 56.36 56.94
UltraEdit 28.12 15.71 7.41 21.15 36.51 35.42 5.48 6.90 5.00 75.38 38.18 43.06
AnyEdit 3.12 10.00 11.11 21.15 25.40 27.08 0.00 1.72 1.67 56.92 40.00 52.78
MagicBrush 7.81 12.86 3.70 19.23 15.87 20.83 0.00 0.00 3.33 43.08 34.55 43.06
IP2P 1.56 8.57 5.56 13.46 15.87 25.00 1.37 0.00 5.00 47.69 20.00 31.94
Table 11: Image rates, overall task rates, and marginal means across three turns (complex mode).
Model Image Success Rate Overall Task Rate Marginal Task Rate
T1 T2 T3 T1 T2 T3 T1 T2 T3
GPT-Image-1 73.08 48.45 28.78 73.08 69.77 68.25 73.08 69.98 70.19
Nano Banana 71.46 46.56 28.14 71.46 68.83 67.27 71.46 69.03 67.21
Gemini 2.0 Flash 67.43 40.63 21.89 67.43 64.54 61.94 67.43 64.80 62.17
Qwen-Image-Edit 72.90 46.15 27.62 72.90 69.23 68.07 72.90 67.83 67.66
Step1X-Edit 61.36 32.34 15.73 61.36 57.69 55.01 61.36 55.77 54.02
FLUX.1-Kontext-dev 60.14 33.74 19.58 60.14 59.53 57.87 60.14 57.52 56.99
OmniGen 54.55 23.43 11.01 54.55 50.96 49.83 54.55 48.95 47.90
AnyEdit 39.86 10.31 2.80 39.86 34.79 34.27 39.86 31.47 31.64
UltraEdit 50.70 22.03 8.22 50.70 48.34 46.62 50.70 47.73 47.38
MagicBrush 42.13 14.86 4.55 42.13 38.46 38.81 42.13 38.64 39.34
IP2P 37.24 12.41 2.80 37.24 37.76 35.14 37.24 38.46 34.62
Table 12: Success rates (%) for five instruction types across three turns (complex mode).
Model Subject Replace Subject Remove Material Alter Color Alter Subject Add
T1 T2 T3 T1 T2 T3 T1 T2 T3 T1 T2 T3 T1 T2 T3
GPT-Image-1 82.22 80.65 78.99 70.31 63.87 58.54 96.49 90.20 84.66 100.00 97.27 91.88 81.67 70.73 69.00
Nano Banana 91.67 88.89 79.33 67.16 55.83 59.88 92.59 85.15 83.33 97.87 98.15 94.19 69.84 71.21 67.77
Gemini 2.0 Flash 85.19 82.88 75.90 57.33 55.56 54.15 93.94 75.00 74.87 100.00 96.90 93.01 70.15 65.97 66.67
Qwen-Image-Edit 87.04 86.49 80.72 70.67 58.33 54.63 93.94 86.21 85.64 100.00 99.22 98.39 77.61 75.69 73.84
Step1X-Edit 90.74 84.68 76.51 52.00 41.67 42.93 93.94 82.76 79.26 100.00 93.02 90.86 64.18 59.03 54.01
FLUX.1-Kontext-dev 85.19 82.88 74.10 54.67 47.92 40.00 86.36 75.00 76.06 100.00 99.22 98.39 67.16 68.75 63.29
OmniGen 88.89 82.88 75.90 46.67 39.58 41.95 86.36 73.28 72.34 100.00 96.12 93.01 53.73 48.61 49.37
AnyEdit 66.67 58.56 49.40 33.33 20.14 22.44 77.27 76.72 70.21 78.57 68.99 72.04 28.36 22.92 21.10
UltraEdit 88.89 87.39 78.31 26.67 30.56 31.71 87.88 79.31 77.66 98.21 93.02 90.32 38.81 46.53 43.46
MagicBrush 83.33 74.77 69.28 28.00 27.78 29.27 83.33 69.83 73.94 92.86 84.50 83.87 37.31 35.42 32.91
IP2P 75.93 73.87 61.45 25.33 31.25 24.88 71.21 63.79 59.57 87.50 82.95 79.03 23.88 31.94 29.11
Table 13: Success rates (%) for four instruction types across three turns (complex mode).
Model Text Change Position Change Count Change Background Change
T1 T2 T3 T1 T2 T3 T1 T2 T3 T1 T2 T3
GPT-Image-1 87.50 92.24 93.75 30.95 22.11 25.36 13.11 10.38 13.12 98.15 96.08 93.37
Nano Banana 84.21 85.34 84.24 24.44 26.00 22.54 20.34 14.15 13.75 98.15 93.40 93.71
Gemini 2.0 Flash 85.94 86.57 84.04 17.31 23.68 16.67 11.11 10.77 6.84 90.77 84.17 80.73
Qwen-Image-Edit 98.44 97.01 93.62 21.15 22.61 24.54 12.33 3.05 3.66 98.46 95.83 93.75
Step1X-Edit 57.81 59.70 52.13 15.38 21.74 25.15 1.37 3.05 1.57 86.15 80.00 73.44
FLUX.1-Kontext-dev 50.00 48.51 47.87 15.38 26.96 27.61 0.00 1.53 3.66 90.77 90.00 88.54
OmniGen 28.12 32.84 27.13 15.38 21.74 17.18 6.85 1.53 2.09 75.38 70.00 69.79
AnyEdit 3.12 5.22 7.98 25.00 26.96 29.45 1.37 1.53 1.57 56.92 44.17 40.62
UltraEdit 29.69 16.42 11.70 21.15 26.09 22.70 5.48 2.29 3.14 75.38 65.00 64.06
MagicBrush 7.81 4.48 6.38 19.23 22.61 17.18 0.00 0.00 5.24 43.08 36.67 35.42
IP2P 3.12 5.22 6.38 13.46 20.00 20.25 1.37 2.29 1.57 47.69 37.50 38.54
Table 14: Consistency scores (%) across DINOv3-based and L1-based object/background metrics. (multi-turn mode)
Model Object DINOv3 Consistency Background DINOv3 Consistency Object $1-L_{1}$ Consistency Background $1-L_{1}$ Consistency
T1 T2 T3 T1 T2 T3 T1 T2 T3 T1 T2 T3
Seedream 4.0 89.50 83.68 81.18 95.52 92.38 90.54 93.31 88.76 86.22 94.30 89.04 85.00
Nano Banana 90.17 85.25 84.38 97.65 95.70 94.58 92.39 90.00 89.10 95.95 94.70 93.88
GPT-Image-1 73.26 68.80 67.21 88.74 86.76 83.78 79.65 78.32 77.60 78.07 76.53 75.79
FLUX.1-Kontext-max 90.91 86.66 83.68 96.95 95.15 93.11 94.16 91.26 89.46 95.88 93.43 91.46
Gemini 2.0 Flash 85.53 77.12 72.02 95.63 93.07 89.74 90.72 86.32 84.11 95.04 93.15 91.73
Qwen-Image-Edit 77.12 71.56 68.51 91.31 89.47 87.45 83.48 79.15 76.32 84.57 81.18 78.43
Step1X-Edit 88.17 81.65 77.33 97.34 95.40 93.09 93.92 90.64 88.80 98.24 97.10 95.73
FLUX.1-Kontext-dev 92.66 87.92 85.29 97.97 96.55 95.14 94.39 91.59 89.91 96.36 95.06 94.13
OmniGen 88.34 80.77 73.64 97.66 96.08 94.21 93.87 91.02 89.43 97.44 97.00 96.34
UltraEdit 78.81 75.11 72.24 94.80 93.89 92.57 91.86 90.47 89.65 97.12 96.62 96.19
AnyEdit 82.02 73.41 63.04 90.82 84.42 77.17 92.52 88.96 84.97 93.72 89.98 86.05
MagicBrush 79.70 70.71 65.46 94.22 91.81 88.27 91.13 87.13 85.56 96.31 94.52 92.87
IP2P 68.24 56.83 48.01 85.47 79.89 72.59 84.44 79.74 77.21 91.31 87.21 83.51
Table 15: Consistency and $1-L_{1}$ metrics across three turns (multi-turn mode for square image).
Model Object DINOv3 Consistency (Mean) Background DINOv3 Consistency Object $L_{1}$ Consistency (Mean) Background $L_{1}$ Consistency
T1 T2 T3 T1 T2 T3 T1 T2 T3 T1 T2 T3
Seedream 4.0 89.04 85.39 80.74 95.76 92.43 88.27 91.66 88.02 84.43 95.34 92.45 89.69
Nano Banana 88.68 85.19 83.31 96.88 93.79 92.41 89.73 87.64 88.48 94.21 93.77 93.36
GPT-Image-1 73.62 71.27 67.27 90.00 85.72 82.16 80.33 78.80 79.88 85.69 81.86 80.85
FLUX.1-Kontext-max 91.34 89.90 86.04 95.99 94.18 90.10 93.81 92.80 90.58 97.27 95.77 93.94
Gemini 2.0 Flash 84.16 79.91 71.13 91.50 91.83 87.07 88.98 87.07 85.13 95.19 93.35 93.34
Qwen-Image-Edit 75.36 72.68 68.85 88.50 88.98 82.71 81.66 78.47 77.04 91.16 87.85 85.78
Step1X-Edit 87.40 84.03 78.27 97.43 92.42 88.15 93.14 91.57 89.40 98.10 97.12 96.01
FLUX.1-Kontext-dev 92.55 88.92 84.04 96.49 93.57 92.19 92.83 90.61 88.34 96.91 95.74 94.73
OmniGen 89.41 84.51 77.77 97.34 93.13 86.64 93.26 91.48 89.53 95.79 96.44 95.36
UltraEdit 79.43 76.51 71.54 92.83 89.88 86.54 91.74 90.16 89.24 95.87 95.14 94.63
AnyEdit 82.25 72.40 53.59 86.12 78.60 70.09 92.33 88.00 82.10 94.36 91.68 87.55
MagicBrush 79.02 75.60 70.07 92.07 87.33 81.08 89.67 87.53 86.90 95.29 93.88 92.38
IP2P 76.38 66.12 54.94 85.29 81.46 65.50 86.90 82.25 77.79 92.45 88.01 84.51
Table 16: Consistency and $L_{1}$ metrics across three turns (multi-turn model for unsquared image).
Model Object DINOv3 Consistency (Mean) Background DINOv3 Consistency Object $1-L_{1}$ Consistency (Mean) Background $1-L_{1}$ Consistency
T1 T2 T3 T1 T2 T3 T1 T2 T3 T1 T2 T3
Seedream 4.0 89.55 83.48 81.23 95.49 92.38 90.86 93.50 88.85 86.44 94.17 88.62 84.35
Nano Banana 90.34 85.26 84.50 97.74 95.94 94.88 92.69 90.29 89.17 96.16 94.81 93.95
GPT-Image-1 73.22 68.53 67.20 88.60 86.88 83.99 79.57 78.27 77.35 77.20 75.91 75.14
FLUX.1-Kontext-max 90.86 86.31 83.43 97.06 95.27 93.51 94.20 91.09 89.34 95.73 93.16 91.14
Gemini 2.0 Flash 85.69 76.79 72.12 96.14 93.23 90.10 90.91 86.23 83.99 95.02 93.12 91.51
Qwen-Image-Edit 77.32 71.43 68.47 91.65 89.53 88.10 83.68 79.23 76.23 83.78 80.37 77.43
Step1X-Edit 88.26 81.37 77.22 97.33 95.77 93.77 94.01 90.53 88.73 98.26 97.09 95.70
FLUX.1-Kontext-dev 92.67 87.80 85.44 98.15 96.92 95.54 94.57 91.70 90.10 96.30 94.98 94.05
OmniGen 88.22 80.33 73.14 97.70 96.45 95.25 93.94 90.97 89.42 97.64 97.07 96.47
UltraEdit 78.74 74.95 72.32 95.04 94.39 93.40 91.87 90.51 89.70 97.26 96.80 96.40
AnyEdit 82.00 73.53 64.18 91.39 85.14 78.14 92.54 89.07 85.31 93.64 89.78 85.84
MagicBrush 79.78 70.14 64.90 94.48 92.37 89.27 91.30 87.08 85.40 96.44 94.59 92.94
IP2P 67.32 55.75 47.18 85.50 79.69 73.57 84.17 79.45 77.14 91.18 87.12 83.38
Table 17: Consistency scores (%) across object/background DINOv3 and $L_{1}$ metrics (complex mode).
Model Object DINOv3 Object $1-L_{1}$ Background DINOv3 Background $1-L_{1}$
T1 T2 T3 T1 T2 T3 T1 T2 T3 T1 T2 T3
GPT-Image-1 73.23 70.02 67.57 79.52 78.03 77.29 88.72 86.77 84.79 77.96 77.38 76.51
Nano Banana 89.23 87.20 86.46 92.15 91.08 90.40 97.39 96.69 95.38 96.39 95.75 95.33
Gemini 2.0 Flash 85.41 80.37 77.38 90.60 88.45 86.53 96.03 94.18 92.83 95.03 94.77 93.46
Qwen-Image-Edit 77.12 76.09 76.69 83.48 83.08 83.11 91.31 91.32 90.51 84.57 84.93 85.35
Step1X-Edit 88.14 85.31 84.38 93.93 92.34 92.11 97.34 96.37 95.44 98.24 98.02 98.04
FLUX.1-Kontext-dev 92.66 90.30 89.19 94.39 92.79 91.61 97.97 96.74 95.40 96.36 95.57 94.04
OmniGen 88.37 85.15 83.14 93.88 92.46 91.06 97.62 97.10 96.07 97.45 97.58 97.40
AnyEdit 81.90 82.94 84.92 92.34 92.43 93.72 90.97 92.63 93.78 94.15 95.11 95.87
UltraEdit 78.81 72.75 71.67 91.86 89.51 88.99 94.80 93.01 92.03 97.12 96.35 96.02
MagicBrush 79.70 75.64 75.53 91.13 89.23 88.69 94.22 94.34 93.13 96.31 96.14 95.83
IP2P 68.24 67.31 69.49 84.44 82.93 83.72 85.47 85.88 86.45 91.31 89.94 90.19
Table 18: Human preference scores, p999, and p99 across three turns (multi-turn mode).
Model Human Preference Score p999 p99
T1 T2 T3 T1 T2 T3 T1 T2 T3
GPT-Image-1 6.6519 6.5898 6.5609 84.73 82.96 82.50 74.38 71.91 70.54
Nano Banana 4.9431 5.1179 5.2638 89.67 90.27 90.56 81.31 82.01 82.09
Gemini 2.0 Flash 4.4386 4.2332 4.0677 90.95 92.21 92.95 83.08 84.79 86.24
Qwen-Image-Edit 5.8591 5.7198 5.1502 89.16 97.20 99.04 79.60 90.51 95.28
Step1X-Edit 4.0577 3.3443 2.7569 92.81 95.39 97.27 85.10 88.46 91.21
FLUX.1-Kontext-dev 5.1192 5.0701 5.0354 90.76 91.94 93.21 82.05 83.03 84.58
OmniGen 4.6099 4.0743 3.4958 93.15 95.24 96.71 85.55 88.24 90.50
AnyEdit 3.6609 2.8017 1.9457 86.70 86.70 86.71 77.54 76.53 75.82
UltraEdit 4.7934 4.6806 4.3598 92.71 94.19 95.82 85.42 86.76 88.34
MagicBrush 3.8465 3.0805 2.3606 91.49 93.33 94.20 83.42 84.70 85.32
IP2P 3.2020 2.3779 1.4418 89.61 91.59 92.44 81.79 83.78 84.77
Table 19: Updated human preference scores, p999 scores, and p99 scores across three turns (complex mode).
Model Human Preference Score p999 p99
T1 T2 T3 T1 T2 T3 T1 T2 T3
GPT-Image-1 6.6328 6.8428 6.9655 85.33 84.14 84.44 74.73 73.92 73.07
Nano Banana 4.9444 5.1700 5.3632 89.65 90.67 91.79 81.02 81.93 82.75
Gemini 2.0 Flash 4.4511 4.5428 4.5732 91.27 92.66 93.48 83.85 85.69 86.75
Qwen-Image-Edit 5.8591 5.8769 5.9155 89.16 90.92 92.23 79.60 81.36 82.62
Step1X-Edit 4.0534 3.9063 3.8648 92.82 93.55 94.05 85.11 85.49 86.01
FLUX.1-Kontext-dev 5.1192 5.2446 5.4645 90.76 91.01 91.29 82.05 81.53 81.55
OmniGen 4.5976 4.3070 3.8122 93.15 93.74 95.65 85.56 86.28 88.57
AnyEdit 3.7020 3.7601 3.8382 86.47 87.16 87.43 77.82 78.60 79.12
UltraEdit 4.7934 4.7647 4.8117 92.71 93.06 93.24 85.42 86.09 86.47
MagicBrush 3.8465 3.6029 3.5523 91.49 91.52 91.64 83.42 83.31 83.07
IP2P 3.2020 3.3552 3.5640 89.61 90.46 90.86 81.79 82.57 82.71
Appendix K Human Agreement
The human study was conducted online through Gradio [^3] https://www.gradio.app/. Annotators were asked to answer a 2-way multiple-choice problem (Yes/No) about an editing instruction, an original image, and an edited image. There were very limited potential participant risks, if they were to be exposed to an image that was disturbing or not safe for work (NSFW). It is because the source images we used were from GEit-Bench Liu et al. (2025), which were not in themselves offensive. Also, our agent already filtered out unsafe images during the first decomposition stage. Furthermore, all edited images from the models were passed through its own NSFW filters which blacked out any potentially unsafe content.
We conducted human study on edits made by four exemplary models—Step1X-Edit, AnyEdit, Gemini-Flash 2.0, and Flux.1-Kontext-dev—on EdiVal-Bench, generated by EdiVal-Agent as described in Section I.2. For each edit, we collected two human ratings, yielding a total of $572\times 4\times 2=4{,}576$ annotations. Depending on the prompt (which affected the editing instruction), each annotation took about 1–2 minutes. Raters were recruited online, each holding at least a bachelor’s degree. They were shown the original image, the edited image, and the corresponding instruction, and were asked a binary question: “Evaluate whether the edited image successfully follows the given instruction.”
Appendix L Counting
Among all subtasks, count_change is the most challenging. Even the best-performing model (GPT-Image-1) achieves a success rate below 25% at turn 1, while most models remain under 5%. We also provide illustrative examples in Figure 15.
(a) Base Image
(b) GPT-Image-1 (✓)
(c) Nano Banana (✓)
(d) Gemini 2.0 Flash (×)
(e) Qwen-Image-Edit (×)
Figure 15: Example of the count_change task: changing the number of paper cups to five.
Appendix M More Quality Examples
Input T1 T2 T3
Seedream 4.0 Refer to caption Refer to caption Refer to caption Refer to caption
Nano Banana Refer to caption Refer to caption Refer to caption Refer to caption
GPT-Image-1 Refer to caption Refer to caption Refer to caption Refer to caption
Gemini 2.0 Flash Refer to caption Refer to caption Refer to caption Refer to caption
FLUX.1-Kontext-max Refer to caption Refer to caption Refer to caption Refer to caption
Qwen-Image-Edit Refer to caption Refer to caption Refer to caption Refer to caption
Figure 16: T1: Change the color of pumpkin to purple; T2: Change the background to forest; T3: Remove fabric orange bow. Row-wise quality examples for the first six models: Seedream 4.0, Nano Banana, GPT-Image-1, Gemini 2.0 Flash, FLUX.1-Kontext-max, and Qwen-Image-Edit. Each row shows generations for Input and three editing turns.
Input T1 T2 T3
Step1X-Edit Refer to caption Refer to caption Refer to caption Refer to caption
FLUX.1-kontext-dev Refer to caption Refer to caption Refer to caption Refer to caption
OmniGen Refer to caption Refer to caption Refer to caption Refer to caption
AnyEdit Refer to caption Refer to caption Refer to caption Refer to caption
UltraEdit Refer to caption Refer to caption Refer to caption Refer to caption
MagicBrush Refer to caption Refer to caption Refer to caption Refer to caption
IP2P Refer to caption Refer to caption Refer to caption Refer to caption
Figure 17: T1: Change the color of pumpkin to purple; T2: Change the background to forest; T3: Remove fabric orange bow. Row-wise quality examples for the remaining models: Step1X-Edit, FLUX.1-kontext-dev, OmniGen, AnyEdit, UltraEdit, MagicBrush, and IP2P.